UVCS Re-commissioning Update

UVCS Re-commissioning

Date Long Contact Activity
Detailed Schedule
Hours  Type  Activities 
October 10, 1998 
DOY 283
DSS-24 16:55-19:45 UT
DSS-16 20:00-01:45 UT
TSTOL and 

UVCS Re-commissioning activities

PHASE I Objectives:     a) Turn on UVCS
                                         b) Achieve temperature stability
                                         c) Activate and check out UVCS flight software

PHASE I Activities:

 1.  Verify reception of Medium and Low-Rate telemetry

 2.  Flight Operations Team powers on UVCS A side electronics

    UVCS power on at 17:27 UT!

 3.  Verify that UVCS is in Initialization Mode and that all
    telemetry points are nominal

   UVCS has been successfully powered on, current
   is nominal, and all temperatures are reasonable as
   of 18:05 UT!

 4.  Verify UVCS receipt of near real time commands

   Have verified that commands have been received
   from UVCS EOF workstation!

 5.  Utilize operations software to verify UVCS memory checksums

    Memory checksum test has verified the integrity
    of UVCS Flight Software!

 6.  Activate main UVCS flight software and verify "Standby" mode

   UVCS switched to "Standby" mode at 19:03 UT!

 7.  Flight Operations Team switches on UVCS substitution heaters
    in the telescope to compensate for absence of detector and mechanism
    power dissipation.

   FOT has switched on substitution heaters at 20%!

 8.  Heater configuration and temperatures are checked

   Endcap heaters turned down to 25% power,
   mirror/occulter mechanism heater turned to
   100%, and visible light detector electronics heater
   turned on to raise its temperature toward nominal

   This completes the Near Real Time (NRT) commanding
   portion of Phase I!

 9.  Telemetry watched for at least 8 hours to establish trends in
    temperatures.  Operational and Substitution Heaters are adjusted
    as necessary to establish safe steady instrument configuration.


FOT has switched substitution heaters to telescope unit to 30%

Evaluation of housekeeping data has determined that all UVCS
instrument telemetry point values including instrument voltages,
currents, temperatures, and mechanism position sensors are within
their nominal ranges!



UVCS Re-commissioning activities

PHASE II Objectives:

Partially open aperture door and re-commission the following mechanisms:

                                       a) Polarizer
                                       b) Mirror/Occulter
                                       c) OVI Grating
                                       d) Ly-alpha Grating
                                       e) Filter Inserters
                                       f) O VI Slit
                                       g) Ly-alpha Slit
                                       h) Roll
                                       i) Pointing

Fully open aperture door
Evaluate alignment and re-orient instrument
Power on O VI and Ly-alpha detectors

PHASE II Activities:

UVCS activities scheduled to begin at 19:05
UT on 11 October 1998 were delayed due

to unavailability of appropriate DSN support.

UVCS Phase II activities began at 21:00 UT
12 October 1998.

 1.    Verify that all temperatures are in nominal range

 2.    Verify or set arming plugs for flight operations

 3.    Check safety flags

 4.    Verify that launch locks are open

 5.    Install/check key tables

 6.    Partially open aperture Door

 7.    Initialize torque motor driven mechanisms

      UVCS commanding resumed at 23:34 UT
      12 October 1998

 8.    Switch on mechanism power

 9.    Re-commission polarizer mechanism

 10.  Re-commission mirror/occulter mechanism

 11.  Re-commission internal occulter mechanism

 12.  Re-commission O VI grating mechanism

      UVCS commanding resumed at 17:54 UT
      13 October 1998

 13.  Re-commission Ly-alpha grating mechanism

       Mechanism positions correctly; Grating stability slightly
       out of specification; Data under review.

 14.  Re-commission filter inserter mechanisms

 15.  Re-commission O VI slit mechanism

 16.  Re-commission Ly-alpha slit mechanism

 17.  Re-commission Roll mechanism

 18.  Re-commission Pointing mechanism

 19.  Open UVCS Aperture Door

 20.  Evaluate UVCS Shadow Sensor Readings

 21.  Establish Alignment of UVCS Occulted Telescope to the Sun

       Alignment at current roll orientation is confirmed
       Alignment at all roll orientations awaits spacecraft
       roll manuever on 16 October 1998

 22.  Roll UVCS 57 degrees counter clockwise to center the field
        of view on the East limb

 23.  Power on O VI detector electronics (low voltage only)

 24.  Verify O VI detector telemetry points

 25.  Power on Ly-alpha detector electronics (low voltage only)

 26.  Verify Ly-alpha detector telemetry points

 27.  FOT to turn off substitution heaters

 28.  Continue to monitor UVCS housekeeping data. Adjust temperatures
        as required

       All housekeeping telemetry points including temperatures
       are nominal!



UVCS Re-commissioning activities

PHASE III Objectives:

Re-commission O VI detector
Ramp up O VI detector high voltage
UVCS re-commissioning first light exposure with O VI detector

Re-commission Visible Light detector

Re-commission Ly-alpha detector
Ramp up Ly-alpha detector high voltage
UVCS re-commissioning first light exposure with Ly-alpha detector

Take radial from 1.5 to 3.0 Solar Radii above East Limb

PHASE III Activities:

 1.    O VI stimulus pulse test

 2.    O VI detector background count rate measurement
        with stimulus pulse and full detector mask

 3.    Insert O VI channel high neutral density filter

 4.     Ramp up O VI detector high voltage

 5.     O VI stimulus pulse and background exposures with
        high voltage ramped up

       Dark count rate is 3.1e-05 counts per second per pixel which is nominal

 6.     Take UV exposure with full O VI detector mask and filter disengaged

       UVCS has taken its "re-commissioning
       first light spectrum!"

      UVCS commanding resumed at 20:12 UT
      14 October 1998

 7.     Load detector mask OVI_237_ANGb in slot 4

 8.    Insert visible light detector neutral density filter

 9.    Visible light detector background count rate measurement

 10.   Disengage filter and take VL exposure

 11.   Ly-alpha stimulus pulse test

 12.   Upload observing program for O VI channel to run during the night

       UVCS O VI Mirror Scan above the East Limb begun
       successfully at 03:53 UT 15 October 1998

       O VI 1032 Angstroms
       1.50 - 2.60 Solar Radii

 13.  Turn off mirror/occulter mechanism (MOM) heater

 14.   Ly-alpha detector background count rate measurement

 15.   Insert Ly-alpha channel high neutral density filter

 16.  Ramp up Ly-alpha detector high voltage

 17.  Ly-alpha background exposures with high voltage
        ramped up

       Dark count rate is 2.5e-05 counts per second per pixel which is nominal

 18.  Ramp down Ly-alpha detector voltage and disengage filter

 19.  Ramp up Ly-alpha detector high voltage

 20.  Take UV exposure with full Ly-alpha detector mask.

       UVCS has taken its Lyman Alpha "re-commissioning
       first light spectrum!"

 21.  Put instrument in safe state in preparation for spacecraft manuvers

 23.  360 degree roll and sun sensor check

 24.  Load and start observing program for 18 October 1998

       This concludes this phase of UVCS re-commissioning!
       Now begin the preparations for November's SPARTAN


Created by Raid Suleiman
Last Modified by Raid Suleiman on 10/09/98
Additional Modification by Rai Wu on 10/20/98